OpenStudy (anonymous):

write the equation of the line containing the point (3 -2) and having a slope of 2 in slope intercept form

OpenStudy (alex_mattucci):

Do you know point slope form? y-y1=m(x-x1)

OpenStudy (alex_mattucci):

substitute in -2 for y1 and 3 for x1

OpenStudy (alex_mattucci):

and substitute in the slope for m.

OpenStudy (alex_mattucci):

you should get y-(-2)=2(x-3)

OpenStudy (alex_mattucci):

Now simplify it! -(-2)= 2

OpenStudy (alex_mattucci):

also distribute 2(the slope) to x and -3

OpenStudy (alex_mattucci):

you should now have y+2=2x-6

OpenStudy (alex_mattucci):

Now solve the rest of the equation. Subtract 2 on both sides. you should get your answer. y=2x-8

OpenStudy (alex_mattucci):